# Evidence-based multimodal fusion on structured EHRs and free-text notes for ICU outcome prediction

### v1

@@ INTRODUCTION
All compared models were implemented using Scikit-learn 30, PyTorch 31, and Hugging Face’s Transformers library 28 in Python 3.8.19. The MLP, ResNet, and FT-Transformer models were built using the original source code on GitHub (https://github.com/yandex-research/rtdl-revisiting-models). The Multimodal approach was modified based on the code on Github (https://github.com/WeiChunLin/Bio_Clinical_BERT_Multimodal_Model).
Model evaluation
For comprehensive model evaluation and comparison, we reported the following metrics: precision, recall, specificity, negative predictive value (NPV), balanced accuracy (BACC), F1 score, area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), Brier score, and negative log-likelihood (NLL). Precision, recall, specificity, and NPV assess model performance for positive or negative samples. In contrast, BACC, F1 score, AUROC, and AUPRC provide a more holistic evaluation of predictive performance. Meanwhile, the Brier score and NLL measure the prediction reliability.
Other fusion settings
In addition to fusing evidence from modalities (structured EHRs and free-text notes), we evaluated our framework under two additional fusion settings: data types and data sources. The illustration is shown in Figure 2.
Figure 2: The illustration of different fusion settings.
Data types: Features can be grouped based on their data types. In our study, we further divided the structured data into two types: numerical and categorical data.
Data Sources: Features can also be grouped according to their sources. In our case, the structured data was split into four sources: demographics, vital signs/lab tests, medical treatments, and comorbidities. Similarly, we categorized the clinical notes into four types: Nursing, Nursing/Other, Physician, and Radiology notes. For each group, we generated evidence using individual encoders and evidential layers. These pieces of evidence were then aggregated through the EF module to produce the final evidence for decision-making.
RESULTS
Data description
@@ RESULTS
Patient chracteristic Mortality PLOS
@@
Model performance
Table 2 presents the comparison of model performance under predictive evaluation metrics (BACC, F1 score, AUROC, and AUPRC) and reliability evaluation metrics (Brier score and NLL) for both mortality prediction (2(a)) and PLOS (2(b)) prediction tasks. The comparison results under precision, recall, specificity, and NPV can be found in Appendix E.
@@
Random Forest x 0.7172$\pm$0.0019 0.3820$\pm$0.0018 0.7988$\pm$0.0010 0.3766$\pm$0.0030 0.1891$\pm$0.0006 0.5639$\pm$0.0015
MLP x 0.7486$\pm$0.0003 0.4043$\pm$0.0019 0.8326$\pm$0.0011 0.4429$\pm$0.0033 0.1677$\pm$0.0015 0.4937$\pm$0.0037
ResNet x 0.7521$\pm$0.0011 0.4113$\pm$0.0007 0.8350$\pm$0.0006 0.4468$\pm$0.0016 0.1654$\pm$0.0006 0.4895$\pm$0.0015
FT-Transformer x 0.7592$\pm$0.0025 0.4166$\pm$0.0021 0.8426$\pm$0.0013 0.4577$\pm$0.0029 0.1648$\pm$0.0014 0.4832$\pm$0.0037
BERT x 0.6300$\pm$0.0014 0.2814$\pm$0.0019 0.6777$\pm$0.0014 0.2037$\pm$0.0012 0.2302$\pm$0.0070 0.6523$\pm$0.0159
BioBERT x 0.6164$\pm$0.0043 0.2759$\pm$0.0021 0.6695$\pm$0.0016 0.2181$\pm$0.0009 0.2214$\pm$0.0021 0.6326$\pm$0.0048
Clinical Bert x 0.6614$\pm$0.0007 0.3080$\pm$0.0009 0.7240$\pm$0.0001 0.2928$\pm$0.0008 0.2132$\pm$0.0030 0.6086$\pm$0.0070
Multimodal (BERT) x x 0.7531$\pm$0.0016 0.4079$\pm$0.0031 0.8398$\pm$0.0006 0.4596$\pm$0.0021 0.1691$\pm$0.0030 0.4992$\pm$0.0077
Multimodal (BioBERT) x x 0.7558$\pm$0.0007 0.4073$\pm$0.0027 0.8362$\pm$0.0006 0.4570$\pm$0.0032 0.1725$\pm$0.0028 0.5084$\pm$0.0088
Multimodal (Clinical Bert) x x 0.7584$\pm$0.0005 0.4218$\pm$0.0019 0.8404$\pm$0.0004 0.4686$\pm$0.0026 0.1606$\pm$0.0022 0.4792$\pm$0.0046
Ours (MLP-BERT) x x 0.7525$\pm$0.0030 0.4466$\pm$0.0042 0.8406$\pm$0.0010 0.4764$\pm$0.0020 0.1345$\pm$0.0050 0.4041$\pm$0.0128
Ours (MLP-BioBERT) x x 0.7492$\pm$0.0042 0.4493$\pm$0.0060 0.8408$\pm$0.0012 0.4758$\pm$0.0009 0.1301$\pm$0.0062 0.3927$\pm$0.0159
Ours (MLP-Clinical Bert) x x 0.7467$\pm$0.0022 0.4629$\pm$0.0033 0.8465$\pm$0.0008 0.4935$\pm$0.0011 0.1176$\pm$0.0027 0.3594$\pm$0.0066
Ours (ResNet-BERT) x x 0.7581$\pm$0.0037 0.4404$\pm$0.0038 0.8415$\pm$0.0011 0.4688$\pm$0.0014 0.1435$\pm$0.0052 0.4296$\pm$0.0133
Ours (ResNet-BioBERT) x x 0.7546$\pm$0.0039 0.4438$\pm$0.0053 0.8424$\pm$0.0012 0.4671$\pm$0.0014 0.1392$\pm$0.0061 0.4187$\pm$0.0157
Ours (ResNet-Clinical Bert) x x 0.7580$\pm$0.0023 0.4472$\pm$0.0042 0.8474$\pm$0.0006 0.4899$\pm$0.0015 0.1387$\pm$0.0043 0.4178$\pm$0.0110
Ours (FT-Trans-BERT) x x 0.7634$\pm$0.0012 0.4432$\pm$0.0060 0.8485$\pm$0.0007 0.4797$\pm$0.0029 0.1450$\pm$0.0049 0.4345$\pm$0.0129
Ours (FT-Trans-BioBERT) x x 0.7610$\pm$0.0015 0.4507$\pm$0.0007 0.8486$\pm$0.0011 0.4796$\pm$0.0027 0.1358$\pm$0.0014 0.4095$\pm$0.0045
Ours (FT-Trans-Clinical Bert) x x 0.7672$\pm$0.0032 0.4541$\pm$0.0032 0.8534$\pm$0.0012 0.4977$\pm$0.0011 0.1390$\pm$0.0043 0.4192$\pm$0.0112
(a) Mortality prediction
@@
Random Forest x 0.6680$\pm$0.0025 0.3492$\pm$0.0038 0.7270$\pm$0.0018 0.2980$\pm$0.0029 0.2350$\pm$0.0009 0.6649$\pm$0.0018
MLP x 0.6845$\pm$0.0009 0.3736$\pm$0.0011 0.7528$\pm$0.0010 0.3353$\pm$0.0010 0.1950$\pm$0.0018 0.5739$\pm$0.0044
ResNet x 0.6851$\pm$0.0017 0.3774$\pm$0.0018 0.7532$\pm$0.0004 0.3366$\pm$0.0021 0.1929$\pm$0.0012 0.5702$\pm$0.0027
FT-Transformer x 0.6956$\pm$0.0011 0.3790$\pm$0.0021 0.7674$\pm$0.0006 0.3404$\pm$0.0021 0.1947$\pm$0.0031 0.5704$\pm$0.0079
BERT x 0.6087$\pm$0.0011 0.2960$\pm$0.0016 0.6507$\pm$0.0015 0.2262$\pm$0.0009 0.2316$\pm$0.0043 0.6557$\pm$0.0091
BioBERT x 0.6057$\pm$0.0016 0.2940$\pm$0.0012 0.6487$\pm$0.0014 0.2214$\pm$0.0012 0.2297$\pm$0.0029 0.6518$\pm$0.0063
Clinical Bert x 0.6420$\pm$0.0016 0.3261$\pm$0.0015 0.6946$\pm$0.0009 0.2661$\pm$0.0006 0.2225$\pm$0.0031 0.6350$\pm$0.0069
Multimodal (BERT) x x 0.6903$\pm$0.0018 0.3739$\pm$0.0012 0.7625$\pm$0.0006 0.3427$\pm$0.0007 0.2017$\pm$0.0028 0.5960$\pm$0.0082
Multimodal (BioBERT) x x 0.6905$\pm$0.0022 0.3731$\pm$0.0025 0.7606$\pm$0.0007 0.3364$\pm$0.0009 0.1974$\pm$0.0026 0.5888$\pm$0.0043
Multimodal (Clinical Bert) x x 0.6933$\pm$0.0014 0.3714$\pm$0.0023 0.7648$\pm$0.0009 0.3544$\pm$0.0018 0.2027$\pm$0.0056 0.5871$\pm$0.0136
Ours (MLP-BERT) x x 0.6878$\pm$0.0016 0.3868$\pm$0.0030 0.7580$\pm$0.0004 0.3499$\pm$0.0015 0.1773$\pm$0.0022 0.5288$\pm$0.0063
Ours (MLP-BioBERT) x x 0.6866$\pm$0.0022 0.3905$\pm$0.0038 0.7573$\pm$0.0010 0.3461$\pm$0.0023 0.1720$\pm$0.0031 0.5172$\pm$0.0076
Ours (MLP-Clinical Bert) x x 0.6870$\pm$0.0012 0.3952$\pm$0.0032 0.7633$\pm$0.0008 0.3557$\pm$0.0021 0.1637$\pm$0.0054 0.4943$\pm$0.0145
Ours (ResNet-BERT) x x 0.6931$\pm$0.0013 0.3887$\pm$0.0040 0.7655$\pm$0.0011 0.3584$\pm$0.0007 0.1815$\pm$0.0045 0.5398$\pm$0.0107
Ours (ResNet-BioBERT) x x 0.6896$\pm$0.0011 0.3918$\pm$0.0037 0.7647$\pm$0.0009 0.3532$\pm$0.0010 0.1752$\pm$0.0048 0.5241$\pm$0.0123
Ours (ResNet-Clinical Bert) x x 0.6974$\pm$0.0011 0.4019$\pm$0.0015 0.7705$\pm$0.0009 0.3639$\pm$0.0010 0.1694$\pm$0.0018 0.5082$\pm$0.0042
Ours (FT-Trans-BERT) x x 0.7005$\pm$0.0010 0.3809$\pm$0.0024 0.7725$\pm$0.0006 0.3524$\pm$0.0011 0.1927$\pm$0.0039 0.5617$\pm$0.0100
Ours (FT-Trans-BioBERT) x x 0.6986$\pm$0.0011 0.3847$\pm$0.0030 0.7701$\pm$0.0012 0.3504$\pm$0.0014 0.1832$\pm$0.0045 0.5376$\pm$0.0113
Ours (FT-Trans-Clinical Bert) x x 0.7027$\pm$0.0011 0.3942$\pm$0.0020 0.7743$\pm$0.0006 0.3575$\pm$0.0014 0.1760$\pm$0.0018 0.5205$\pm$0.0038
(b) PLOS prediction
Table 2: The model comparison for mortality (a) and PLOS prediction (b) tasks.
@@
Overall, our framework demonstrated superior performance across both tasks. In the mortality prediction task, our framework using MLP and Clinical BERT as the backbone achieved the highest F1 score (0.4629) and the lowest Brier score (0.1176) and NLL (0.3594). Furthermore, with FT-Transformer and Clinical BERT, it achieved the highest balanced accuracy (BACC) of 0.7672, AUROC of 0.8534, and AUPRC of 0.4977. Compared to the best baseline models, our framework improved predictive performance by approximately 1.05% in BACC, 9.74% in F1 score, 1.28% in AUROC, and 6.21% in AUPRC. Additionally, it significantly enhanced prediction reliability, with reductions of about 26.8% in Brier score and 25.0% in NLL.
In the PLOS prediction task, our framework demonstrated exceptional performance across all key predictive evaluation metrics, achieving a BACC of 0.7027, an F1 score of 0.4019, an AUROC of 0.7743, and an AUPRC of 0.3639. Compared to the best baseline model, our framework showed notable improvements: a 1.02% increase in BACC, a 6.04% increase in F1 score, a 0.9% increase in AUROC, and a 2.68% increase in AUPRC. Additionally, it enhanced prediction reliability with a 15.1% reduction in Brier score and a 13.3% reduction in NLL.
Models that relied solely on structured data or free-text notes underperformed compared to our proposed framework in both predictive and reliability evaluation metrics. While the multimodal approach showed marginally better performance than other baselines in the mortality prediction task, it delivered weaker results in the PLOS prediction task.
Evaluation on different fusion settings
Figure 3 illustrates the evaluation of our framework across three fusion settings applied to different backbone models for mortality prediction. For the corresponding results in the PLOS prediction task, please refer to Appendix F. Both tasks exhibited similar patterns in the analysis.

### v2

@@ 4.5 Implementation details
Each model was trained five times with different random seeds, and we reported the average results along with the standard error of the metrics to ensure statistical reliability. Hyperparameters were optimized for all baseline models to achieve the best results. All compared models were implemented using Scikit-learn [44], PyTorch [45], and Hugging Face’s Transformers library [42] in Python 3.8.19. The MLP, ResNet, and FT-Transformer models were built using the original source code on GitHub  [^3] https://github.com/yandex-research/rtdl-revisiting-models.. The Multimodal approach was modified based on the code on Github  [^4] https://github.com/WeiChunLin/Bio_Clinical_BERT_Multimodal_Model..
4.6 Model evaluation
For comprehensive model evaluation and comparison, we reported the two types of metrics: predictive accuracy and reliability.
  • Predictive accuracy ensures that models correctly identify critically ill patients who require urgent intervention, thereby reducing the risk of misdiagnosis and unnecessary treatments. To comprehensively assess accuracy, we consider two aspects:
    • Class-specific accuracy metrics which evaluate performance separately for positive and negative cases using metrics such as precision, recall, specificity, and negative predictive value (NPV).
      Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}},$ (13)
      Recall $=\frac{\text{TP}}{\text{TP}+\text{FN}},$ (14)
      Specificity $=\frac{\text{TN}}{\text{TN}+\text{FP}},$ (15)
      NPV $=\frac{\text{TN}}{\text{TN}+\text{FN}},$ (16)
      where TP, TN, FP, and FN are True Positive, True Negative, False Positive, and False Negative, respectively.
    • Holistic accuracy metrics include balanced accuracy (BACC), F1 score, the area under the receiver operating characteristic curve (AUROC), and the area under the precision-recall curve (AUPRC). AUROC is determined by calculating the area under the ROC curve (TP rate against FP rate across different threshold settings) and AUPRC calculates the area under the Precision-Recall curve across various threshold settings.
      BACC $=\frac{1}{2}\left(\text{Recall}+\text{Specificity}\right),$ (17)
      F1 $=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}.$ (18)
      BACC, F1 score, and AUPRC are well-suited for evaluating model performance in imbalanced class distributions, as they provide a more balanced assessment than traditional accuracy measures.
  • Reliability metrics quantify the model’s confidence in predictions. A model with high reliability provides well-calibrated probability estimates, allowing clinicians to make informed risk assessments. Here, we evaluate the performance of the Brier score and negative log-likelihood (NLL).
    Brier Score $=\frac{1}{N}\sum_{i=1}^{N}(p_{i}-o_{i})^{2},$ (19)
    NLL $=-\frac{1}{N}\sum_{i=1}^{N}\left(o_{i}\log(p_{i})+(1-o_{i})\log(1-p_{i})\right),$ (20)
    where $N$ is the number of instances, $p_{i}$ is the predicted probability of the positive class for instance $i$, and $o_{i}$ is the actual outcome for instance $i$ (1 if positive, 0 if negative).
5 RESULTS
5.1 Data description
@@
Patient characteristic Mortality PLOS
@@
5.2 Model performance
5.2.1 Predictive accuracy
Table 2 and 3 present the comparison of model performance for holistic predictive accuracy in mortality and PLOS prediction tasks, respectively.
@@ 5.2.1 Predictive accuracy
Random Forest x 0.7172$\pm$0.0019 0.3820$\pm$0.0018 0.7988$\pm$0.0010 0.3766$\pm$0.0030
MLP x 0.7486$\pm$0.0003 0.4043$\pm$0.0019 0.8326$\pm$0.0011 0.4429$\pm$0.0033
ResNet x 0.7521$\pm$0.0011 0.4113$\pm$0.0007 0.8350$\pm$0.0006 0.4468$\pm$0.0016
FT-Transformer x 0.7592$\pm$0.0025 0.4166$\pm$0.0021 0.8426$\pm$0.0013 0.4577$\pm$0.0029
BERT as text encoder
Text encoder only x 0.6300$\pm$0.0014 0.2814$\pm$0.0019 0.6777$\pm$0.0014 0.2037$\pm$0.0012
Multimodal x x 0.7531$\pm$0.0016 0.4079$\pm$0.0031 0.8398$\pm$0.0006 0.4596$\pm$0.0021
@@
Ours (ResNet) x x 0.7581$\pm$0.0037 0.4404$\pm$0.0038 0.8415$\pm$0.0011 0.4688$\pm$0.0014
Ours (FT-Trans) x x 0.7634$\pm$0.0012 0.4432$\pm$0.0060 0.8485$\pm$0.0007 0.4797$\pm$0.0029
BioBERT as text encoder
Text encoder only x 0.6164$\pm$0.0043 0.2759$\pm$0.0021 0.6695$\pm$0.0016 0.2181$\pm$0.0009
Multimodal x x 0.7558$\pm$0.0007 0.4073$\pm$0.0027 0.8362$\pm$0.0006 0.4570$\pm$0.0032
[…]
Ours (ResNet) x x 0.7546$\pm$0.0039 0.4438$\pm$0.0053 0.8424$\pm$0.0012 0.4671$\pm$0.0014
Ours (FT-Trans) x x 0.7610$\pm$0.0015 0.4507$\pm$0.0007 0.8486$\pm$0.0011 0.4796$\pm$0.0027
Clinical BERT as text encoder
Text encoder only x 0.6614$\pm$0.0007 0.3080$\pm$0.0009 0.7240$\pm$0.0001 0.2928$\pm$0.0008
Multimodal x x 0.7584$\pm$0.0005 0.4218$\pm$0.0019 0.8404$\pm$0.0004 0.4686$\pm$0.0026
[…]
Ours (ResNet) x x 0.7580$\pm$0.0023 0.4472$\pm$0.0042 0.8474$\pm$0.0006 0.4899$\pm$0.0015
Ours (FT-Trans) x x 0.7672$\pm$0.0032 0.4541$\pm$0.0032 0.8534$\pm$0.0012 0.4977$\pm$0.0011
Table 2: Comparison of predictive accuracy on mortality prediction. The best results among models using the same text encoder are in bold, and the overall best results are shaded in grey.
[…]
Random Forest x 0.6680$\pm$0.0025 0.3492$\pm$0.0038 0.7270$\pm$0.0018 0.2980$\pm$0.0029
MLP x 0.6845$\pm$0.0009 0.3736$\pm$0.0011 0.7528$\pm$0.0010 0.3353$\pm$0.0010
ResNet x 0.6851$\pm$0.0017 0.3774$\pm$0.0018 0.7532$\pm$0.0004 0.3366$\pm$0.0021
FT-Transformer x 0.6956$\pm$0.0011 0.3790$\pm$0.0021 0.7674$\pm$0.0006 0.3404$\pm$0.0021
BERT as text encoder
Text encoder only x 0.6087$\pm$0.0011 0.2960$\pm$0.0016 0.6507$\pm$0.0015 0.2262$\pm$0.0009
Multimodal x x 0.6903$\pm$0.0018 0.3739$\pm$0.0012 0.7625$\pm$0.0006 0.3427$\pm$0.0007
[…]
Ours (ResNet) x x 0.6931$\pm$0.0013 0.3887$\pm$0.0040 0.7655$\pm$0.0011 0.3584$\pm$0.0007
Ours (FT-Trans) x x 0.7005$\pm$0.0010 0.3809$\pm$0.0024 0.7725$\pm$0.0006 0.3524$\pm$0.0011
BioBERT as text encoder
Text encoder only x 0.6057$\pm$0.0016 0.2940$\pm$0.0012 0.6487$\pm$0.0014 0.2214$\pm$0.0012
Multimodal x x 0.6905$\pm$0.0022 0.3731$\pm$0.0025 0.7606$\pm$0.0007 0.3364$\pm$0.0009
[…]
Ours (ResNet) x x 0.6896$\pm$0.0011 0.3918$\pm$0.0037 0.7647$\pm$0.0009 0.3532$\pm$0.0010
Ours (FT-Trans) x x 0.6986$\pm$0.0011 0.3847$\pm$0.0030 0.7701$\pm$0.0012 0.3504$\pm$0.0014
Clinical BERT as text encoder
Text encoder only x 0.6420$\pm$0.0016 0.3261$\pm$0.0015 0.6946$\pm$0.0009 0.2661$\pm$0.0006
Multimodal x x 0.6933$\pm$0.0014 0.3714$\pm$0.0023 0.7648$\pm$0.0009 0.3544$\pm$0.0018
[…]
Ours (ResNet) x x 0.6974$\pm$0.0011 0.4019$\pm$0.0015 0.7705$\pm$0.0009 0.3639$\pm$0.0010
Ours (FT-Trans) x x 0.7027$\pm$0.0011 0.3942$\pm$0.0020 0.7743$\pm$0.0006 0.3575$\pm$0.0014
Table 3: Comparison of predictive accuracy on PLOS prediction. The best results among models using the same text encoder are in bold, and the overall best results are shaded in grey.
[…]
Random Forest x 0.1891$\pm$0.0006 0.5639$\pm$0.0015
MLP x 0.1677$\pm$0.0015 0.4937$\pm$0.0037
ResNet x 0.1654$\pm$0.0006 0.4895$\pm$0.0015
FT-Transformer x 0.1648$\pm$0.0014 0.4832$\pm$0.0037
BERT as text encoder
Text encoder only x 0.2302$\pm$0.0070 0.6523$\pm$0.0159
Multimodal x x 0.1691$\pm$0.0030 0.4992$\pm$0.0077
[…]
Ours (ResNet) x x 0.1435$\pm$0.0052 0.4296$\pm$0.0133
Ours (FT-Trans) x x 0.1450$\pm$0.0049 0.4345$\pm$0.0129
BioBERT as text encoder
Text encoder only x 0.2214$\pm$0.0021 0.6326$\pm$0.0048
Multimodal x x 0.1725$\pm$0.0028 0.5084$\pm$0.0088
[…]
Ours (ResNet) x x 0.1392$\pm$0.0061 0.4187$\pm$0.0157
Ours (FT-Trans) x x 0.1358$\pm$0.0014 0.4095$\pm$0.0045
Clinical BERT as text encoder
Text encoder only x 0.2132$\pm$0.0030 0.6086$\pm$0.0070
Multimodal x x 0.1606$\pm$0.0022 0.4792$\pm$0.0046
[…]
Ours (ResNet) x x 0.1387$\pm$0.0043 0.4178$\pm$0.0110
Ours (FT-Trans) x x 0.1390$\pm$0.0043 0.4192$\pm$0.0112
Table 4: Comparison of reliability performance on mortality prediction.
[…]
Random Forest x 0.2350$\pm$0.0009 0.6649$\pm$0.0018
MLP x 0.1950$\pm$0.0018 0.5739$\pm$0.0044
ResNet x 0.1929$\pm$0.0012 0.5702$\pm$0.0027
FT-Transformer x 0.1947$\pm$0.0031 0.5704$\pm$0.0079
BERT as text encoder
Text encoder only x 0.2316$\pm$0.0043 0.6557$\pm$0.0091
Multimodal x x 0.2017$\pm$0.0028 0.5960$\pm$0.0082
[…]
Ours (ResNet) x x 0.1815$\pm$0.0045 0.5398$\pm$0.0107
Ours (FT-Trans) x x 0.1927$\pm$0.0039 0.5617$\pm$0.0100
BioBERT as text encoder
Text encoder only x 0.2297$\pm$0.0029 0.6518$\pm$0.0063
Multimodal x x 0.1974$\pm$0.0026 0.5888$\pm$0.0043
[…]
Ours (ResNet) x x 0.1752$\pm$0.0048 0.5241$\pm$0.0123
Ours (FT-Trans) x x 0.1832$\pm$0.0045 0.5376$\pm$0.0113
Clinical BERT as text encoder
Text encoder only x 0.2225$\pm$0.0031 0.6350$\pm$0.0069
Multimodal x x 0.2027$\pm$0.0056 0.5871$\pm$0.0136
[…]
Ours (ResNet) x x 0.1694$\pm$0.0018 0.5082$\pm$0.0042
Ours (FT-Trans) x x 0.1760$\pm$0.0018 0.5205$\pm$0.0038
Table 5: Comparison of reliability performance on PLOS prediction.
@@
Overall, our framework demonstrated superior performance across both tasks. In the mortality prediction task, our framework using MLP and Clinical BERT as the backbones achieved the highest F1 score (0.4629). Furthermore, with FT-Transformer and Clinical BERT, it achieved the highest BACC of 0.7672, AUROC of 0.8534, and AUPRC of 0.4977. Compared to the best baseline models, our framework improved predictive performance by approximately 1.05% in BACC, 9.74% in F1 score, 1.28% in AUROC, and 6.21% in AUPRC. In the PLOS prediction task, our framework demonstrated exceptional performance across all evaluation metrics for predictive accuracy , achieving a BACC of 0.7027, an F1 score of 0.4019, an AUROC of 0.7743, and an AUPRC of 0.3639. Compared to the best baselines, our framework showed notable improvements: a 1.02% increase in BACC, a 6.04% increase in F1 score, a 0.9% increase in AUROC, and a 2.68% increase in AUPRC.
Over the baseline models, the multimodal approaches showed marginally worse predictive accuracy than others in both tasks. Specifically, in the mortality prediction task, the multimodal approach with clinical BERT as the backbone achieved the highest F1 of 0.4218 and AUPRC of 0.4686. In the PLOS task, it achieved the highest AUPRC of 0.3544.
5.2.2 Reliability
In clinical settings, assessing prediction reliability is as important as evaluating predictability. Table 4 and 5 present the comparison of model performance based on reliability across both prediction tasks.
In the mortality prediction task, our framework achieved the lowest Brier score (0.1176) and NLL (0.3594) with MLP and Clinical BERT as the backbones. It improved prediction reliability significantly compared to the best baseline models, reducing the Brier score by about 26.8% and NLL by 25.0%. In the PLOS prediction task, our framework also showed strong performance, with a Brier score of 0.1637 and an NLL of 0.4943. These results showed notable improvements over the best baseline model, with a 15.1% reduction in Brier score and a 13.3% reduction in NLL.
Compared to other baselines, the multimodal approaches exhibited slightly better reliability in the mortality prediction task, achieving a Brier score of 0.1606 and an NLL of 0.4792. However, in the PLOS prediction tasks, the multimodal approaches demonstrated weaker reliability.
Figure 3: The illustration of different fusion settings.
5.3 Evaluation on different fusion settings
We also validate the effectiveness of our fusion setting in EHR groups by comparing it with two other additional fusion settings: data types and data sources. The illustration is shown in Figure 3. For the data types fusion, we divided the structured data into two types: numerical and categorical data. For the data sources fusion, we split the structured data into four sources: demographics, vital signs/lab tests, medical treatments, and comorbidities, and categorized the clinical notes into four types: Nursing, Nursing/Other, Physician, and Radiology notes. Figures 4 and 5 illustrate the evaluation of our framework using clinical BERT as text encoder across three fusion settings for both tasks. Evaluation results about models using BERT and BioBERT as text encoders can be found in Appendix A.